# A generalization of Frenkel’s formula

### v1

@@ 1. Introduction
We show the following equivalent formula analogous to [12, Proposition 4.2] and [4, eq. (10)]:
@@ 2. Proof of (7)
To prove (7) we follow the proof of Frenkel’s integral formula in [12, Section 4.1]. Recall [12, eq. (77)] and [3, Theorem B.1] for $A,B>0$:
@@ 3. An infinite dimensional case
It is stated in [10, Section 2.2] that (6) holds for $A,B\in\mathcal{S}_{1,+}$ as follows. Assume first that $\operatorname{supp}A\not\subseteq\operatorname{supp}B$. Then as in Theorem 1 both sides of (1) are $\infty$. Assume second that $\operatorname{supp}A\subseteq\operatorname{supp}B$. As in Theorem 1 we can restrict $A$ and $B$ to $\operatorname{range}B$. Thus, it is enough to discuss the case where $B>0$. Thus, we can summarize [10, Theorem 1] in this form:
@@
Recall Rellich’s theorem [9, Theorem 3.9]. The eigenvalues and the eigenvectors of the pencil $L(\gamma)$ are analytic on $\mathbb{R}$. Hence each nonzero analytic eigenvalue of $L(\gamma)$ has a countable number of zeros that accumulate to $\pm\infty$. Therefore $\|L_{+}(\gamma)\|_{p}$ is a continuous function on $\mathbb{R}$. (See Theorem 4 for the case where $L_{0},L_{1}$ have finite dimensional range.) Define
@@ Acknowledgment
The author thanks Gilad Gour for pointing out the references [4, 5].
@@ A Hermitian pencils is defined:
(2) Assume that $H(z)$ is $1$-indecomposable singular Hermitian pencil. Then a variation of the Kronecker canonical form for general pencils gives decomposition (27) [14, Theorem 1] or [11, Theorem 6.1]. Next observe that for $H_{j}(z)$ has exactly on eigenvector corresponding 0 eigenvalue. As $H_{j}(z)$ is a real symmetric matrix it for a real $z$ it follows that 0-eigenvalue is of multiplicity one of $H_{j}(z)$. Hence, 0 is an eigenvalue of multiplicity $k$ in $\oplus_{j=1}^{k+1}H_{j}(z)$. Therefore, $\det(\lambda I_{n}-H(z))=\lambda^{k}p(\lambda,z)$ where the joint degree of $p(\lambda,z)$ is nkn-k. Furhtermore $p(0,z)$ is a nonzero polynomial of degree at most nkn-k. (if $p(0,z)$ vanish identically, we will deduce that 0 is an eigenvalue of multiplicity at $k+1$.) Hence the product of the nonvanishing nkn-k eigenvalues of $H(z)$ vanish at the zeros of $p(0,z)=0$, which is a set of cardinality at most nkn-k. ∎
@@ A.2. An infinite dimensional case
Recall that if the sequences $\langle\mathbf{y},T_{n}\mathbf{x}\rangle$ ($\{T_{n}\mathbf{x}\}$) converge for each $\mathbf{x},\mathbf{y}\in\mathcal{H}$ ($\mathbf{x}\in\mathcal{H}$) then $\sup_{n\in\mathbb{N}}\|T_{n}\|<\infty$. Furthermore, there exists $T\in\mathcal{B}$ such that $\lim T_{n}\stackrel{{\scriptstyle weak}}{{\to}}T$ ($\lim T_{n}\stackrel{{\scriptstyle strong}}{{\to}}T$) [13, Sec. VI.1].
Let $T\in\mathcal{S}$. One associates with $T$ a family of projective-valued measures on the interval $\|T\|,\|T\|]$, see [13, VIII.3,Proposition]. Then, for any bounded Borel function ffon $\|T\|,\|T\|]$ one can define $f(T)=\int_{-\|T\|}^{\|T\|}f(\lambda)dP_{\lambda}$ such that $\langle\mathbf{x},f(T)\mathbf{x}\rangle=\int_{-\|T\|}^{\|T\|}f(\lambda)d\langle\mathbf{x},P_{\lambda}\mathbf{x}\rangle$. In particular
@@
$\log B=\int_{0}^{\infty}\big((Id+x\mathbf{1})^{-1}-(B+x\mathbf{1})^{-1})dx,\quad B\in\mathcal{S}^{\prime}_{++}.$
@@ A.3. Approximation lemmas
(3) Clearly, $P_{n}\stackrel{{\scriptstyle strong}}{{\to}}\mathbf{1}$. Recall [13, Sec. VI, Problem (6d)] that if $A_{n},B_{n}\in\mathcal{S}$ converge in the strong topology to $A,B\in\mathrm{S}(\mathcal{B})$ respectively, then $A_{n}B_{n}\stackrel{{\scriptstyle strong}}{{\to}}AB$. Hence
@@
Problem 14(b) in [13, Sec. VI] states that if $A_{n}\in\mathrm{S}_{+}(\mathcal{B})$ converges strongly to $A\in\mathrm{S}_{+}(\mathcal{B})$ then $\sqrt{A_{n}}\stackrel{{\scriptstyle strong}}{{\to}}\sqrt{A}$. Hence

### v2

@@ 1. Introduction
We show the following equivalent formula analogous to [8, eq. (1.6)],[13, Proposition 4.2] and [4, eq. (10)]:
@@ 2. Proof of (7)
To prove (7) we follow the proof of Frenkel’s integral formula in [13, Section 4.1]. Recall [13, eq. (77)] and [3, Theorem B.1] for $A,B>0$:
@@ 3. An infinite dimensional case
It is stated in [11, Section 2.2] that (6) holds for $A,B\in\mathcal{S}_{1,+}$ as follows. Assume first that $\operatorname{supp}A\not\subseteq\operatorname{supp}B$. Then as in Theorem 1 both sides of (1) are $\infty$. Assume second that $\operatorname{supp}A\subseteq\operatorname{supp}B$. As in Theorem 1 we can restrict $A$ and $B$ to $\operatorname{range}B$. Thus, it is enough to discuss the case where $B>0$. Thus, we can summarize [11, Theorem 1] in this form:
@@
Recall Rellich’s theorem [10, Theorem 3.9]. The eigenvalues and the eigenvectors of the pencil $L(\gamma)$ are analytic on $\mathbb{R}$. Hence each nonzero analytic eigenvalue of $L(\gamma)$ has a countable number of zeros that accumulate to $\pm\infty$. Therefore $\|L_{+}(\gamma)\|_{p}$ is a continuous function on $\mathbb{R}$. (See Theorem 4 for the case where $L_{0},L_{1}$ have finite dimensional range.) Define
@@
Problem 1.
  1. Assume that $A,B\in\mathcal{S}_{+}$ and $A^{2}\leqslant\alpha^{2}B^{2}$. Then (46) holds. Loewner’s theorem [10, Theorem 4.12] yields $A\leqslant\alpha B$.
  2. (1)
    Is it true that the sequence $D_{n},n\in\mathbb{N}$ in Theorem 3 converges to $BD[B](A)$ in the strong topology in case (1), and in norm in case (2)?
  3. (2)
    Assume that the assumption of part (2) of Lemma 4 holds. Is it true that the sequence $C_{n},n\in\mathbb{N}$ in Theorem 3 converges to $A(\log A-\log B)$ in the strong topology in case (1), and in norm in case (2)?
@@ Acknowledgment
The author thanks Gilad Gour for pointing out the references [4, 5], and Marco Tomamichel for the reference [8].
@@ A Hermitian pencils is defined:
(2) Assume that $H(z)$ is $1$-indecomposable singular Hermitian pencil. Then a variation of the Kronecker canonical form for general pencils gives decomposition (27) [15, Theorem 1] or [12, Theorem 6.1]. Next observe that for $H_{j}(z)$ has exactly on eigenvector corresponding 0 eigenvalue. As $H_{j}(z)$ is a real symmetric matrix it for a real $z$ it follows that 0-eigenvalue is of multiplicity one of $H_{j}(z)$. Hence, 0 is an eigenvalue of multiplicity $k$ in $\oplus_{j=1}^{k+1}H_{j}(z)$. Therefore, $\det(\lambda I_{n}-H(z))=\lambda^{k}p(\lambda,z)$ where the joint degree of $p(\lambda,z)$ is nkn-k. Furhtermore $p(0,z)$ is a nonzero polynomial of degree at most nkn-k. (if $p(0,z)$ vanish identically, we will deduce that 0 is an eigenvalue of multiplicity at $k+1$.) Hence the product of the nonvanishing nkn-k eigenvalues of $H(z)$ vanish at the zeros of $p(0,z)=0$, which is a set of cardinality at most nkn-k. ∎
@@ A.2. An infinite dimensional case
Recall that if the sequences $\langle\mathbf{y},T_{n}\mathbf{x}\rangle$ ($\{T_{n}\mathbf{x}\}$) converge for each $\mathbf{x},\mathbf{y}\in\mathcal{H}$ ($\mathbf{x}\in\mathcal{H}$) then $\sup_{n\in\mathbb{N}}\|T_{n}\|<\infty$. Furthermore, there exists $T\in\mathcal{B}$ such that $\lim T_{n}\stackrel{{\scriptstyle weak}}{{\to}}T$ ($\lim T_{n}\stackrel{{\scriptstyle strong}}{{\to}}T$) [14, Sec. VI.1].
Let $T\in\mathcal{S}$. One associates with $T$ a family of projective-valued measures on the interval $\|T\|,\|T\|]$, see [14, VIII.3,Proposition]. Then, for any bounded Borel function ffon $\|T\|,\|T\|]$ one can define $f(T)=\int_{-\|T\|}^{\|T\|}f(\lambda)dP_{\lambda}$ such that $\langle\mathbf{x},f(T)\mathbf{x}\rangle=\int_{-\|T\|}^{\|T\|}f(\lambda)d\langle\mathbf{x},P_{\lambda}\mathbf{x}\rangle$. In particular
@@
(30) $\log B=\int_{0}^{\infty}\big((Id+x\mathbf{1})^{-1}-(B+x\mathbf{1})^{-1})dx,\quad B\in\mathcal{S}^{\prime}_{++}.$
@@ A.3. Approximation lemmas
(3) Clearly, $P_{n}\stackrel{{\scriptstyle strong}}{{\to}}\mathbf{1}$. Recall [14, Sec. VI, Problem (6d)] that if $A_{n},B_{n}\in\mathcal{S}$ converge in the strong topology to $A,B\in\mathrm{S}(\mathcal{B})$ respectively, then $A_{n}B_{n}\stackrel{{\scriptstyle strong}}{{\to}}AB$. Hence
@@
Problem 14(b) in [14, Sec. VI] states that if $A_{n}\in\mathrm{S}_{+}(\mathcal{B})$ converges strongly to $A\in\mathrm{S}_{+}(\mathcal{B})$ then $\sqrt{A_{n}}\stackrel{{\scriptstyle strong}}{{\to}}\sqrt{A}$. Hence
@@
A.4. Integral representation of two operators
In this subsection we give two convergent formulas for the operators appearing in $\Delta(A\|B)$:
Lemma 4.
Assume that $A\in\mathcal{S},B\in\mathcal{S}_{+}$, and $A^{2}\leqslant\alpha^{2}B^{2}$ for some $\alpha>0$.
  1. (1)
    The following integral formula for $BD[B](A)$ converges, and its norm is bounded from above:
    (46) $BD[B](A)=\int_{0}^{\infty}B(B+x\mathbf{1})^{-1}A(B+x\mathbf{1})^{-1}dx,$
    $\|BD[B](A)\|\leqslant\alpha\|B\|.$
  2. (2)
    Suppose that $A\in\mathcal{S}_{+}$. Then the following integral formula converges
    (47) $A(\log A-\log B)=\int_{0}^{\infty}\big(A(B+x\mathbf{1})^{-1}-A(A+x\mathbf{1})^{-1}\big)dx$
    if one of the following conditions hold:
    1. (a)
      The inequality $B^{2}\leqslant\beta^{2}A^{2}$ holds for some $\beta\geqslant\frac{1}{\alpha}$. Then
      (48) $\|A(\log A-\log B)\|\leqslant\alpha(1+\beta)\|A\|\|B\|\frac{\log\|B\|-\log\|A\|}{\|B\|-\|A\|}.$
    2. (b)
      The inequality $(A-B)^{2}\leqslant\beta^{2}A^{2}$ holds. Then
      (49) $\|A(\log A-\log B)\|\leqslant\alpha\beta\|A\|\|B\|\frac{\log\|B\|-\log\|A\|}{\|B\|-\|A\|}.$
Proof.
Since $A^{2}\leqslant\alpha^{2}B^{2}$, we deduce that $\ker B\subseteq\ker A$. Hence, without loss of generality we can assume that $B\in\mathcal{S}_{++}$. For $\varepsilon\in[0,1]$ let $A_{\varepsilon}=A+\varepsilon\mathbf{1},B_{\varepsilon}=A+\varepsilon\mathbf{1}$.
(1) Fix $x>0$, and $\varepsilon\in(0,1]$. Clearly, $A^{2}\leqslant\alpha^{2}B^{2}_{\varepsilon}$. The equality (31) yields:
$BD[B_{\varepsilon}](A)=\int_{0}^{\infty}B(B+(x+\varepsilon)\mathbf{1})^{-1}A(B+(x+\varepsilon)\mathbf{1})^{-1}dx.$
Observe that for $x>0$
$\|B(B+x\mathbf{1})^{-1}A(B+x\mathbf{1})^{-1}\|\leqslant\|B(B+x\mathbf{1})^{-1}\|\|A(B+x\mathbf{1})^{-1}\|\leqslant$
$\|B(B+x\mathbf{1})^{-1}\|\|(B+x\mathbf{1})^{-1}A^{2}(B+x\mathbf{1})^{-1}\|^{1/2},$
$(B+x\mathbf{1})^{-1}A^{2}(B+x\mathbf{1})^{-1}\leqslant(B+x\mathbf{1})^{-1}\alpha^{2}B^{2}(B+x\mathbf{1})^{-1}\Rightarrow$
$\|(B+x\mathbf{1})^{-1}A^{2}(B+x\mathbf{1})^{-1}\|\leqslant\alpha^{2}\|(B+x\mathbf{1})^{-1}B^{2}(B+x\mathbf{1})^{-1}\|=\alpha^{2}\|B(B+x\mathbf{1})^{-1}\|^{2},$
$\|B(B+x\mathbf{1})^{-1}\|\leqslant\max_{t\in[0,\|B\|]}\frac{t}{t+x}=\frac{\|B\|}{\|B\|+x}.$
(Observe that this inequality is equality.) Hence,
$\|B(B+x\mathbf{1})^{-1}A(B+x\mathbf{1})^{-1}\|\leqslant\frac{\alpha\|B\|^{2}}{(\|B\|+x)^{2}}\Rightarrow$
$\|BD[B_{\varepsilon}](A)\|\leqslant\int_{0}^{\infty}\|B(B+(x+\varepsilon)\mathbf{1})^{-1}A(B+(x+\varepsilon)\mathbf{1})^{-1}\|dx\leqslant$
$\int_{0}^{\infty}\frac{\alpha\|B\|^{2}}{(\|B\|+x+\varepsilon))^{2}}dx=\alpha\frac{\|B\|^{2}}{\|B\|+\varepsilon}<\alpha\|B\|.$
These inequalities shows that the integral in (46) converges. The Lebesgue’s dominated convergence theorem yields that for $\varepsilon_{n}=1/n,n\in\mathbb{N}$ the integrals for $BD[B_{1/n}](A)$ converge in the norm $\|\cdot\|$ to the integral in (46). (See the proof of Theorem 3.) This proves (46).
(2) Use (30) to deduce
$A(\log A_{\varepsilon}-\log B_{\varepsilon})=\int_{0}^{\infty}A\big((B_{\varepsilon}+x\mathbf{1})^{-1}-(A_{\varepsilon}+x\mathbf{1})^{-1}\big)dx=$
$\int_{0}^{\infty}A\left(B_{\varepsilon}+x\mathbf{1})^{-1}\big((A_{\varepsilon}+x\mathbf{1})-(B_{\varepsilon}+x\mathbf{1})(A_{\varepsilon}+x\mathbf{1})^{-1}\right)dx=$
$\int_{0}^{\infty}A(B_{\varepsilon}+x\mathbf{1})^{-1}(A-B)(A_{\varepsilon}+x\mathbf{1})^{-1}dx=F(\varepsilon)-G(\varepsilon),$
$F(\varepsilon)=\int_{0}^{\infty}A(B_{\varepsilon}+x\mathbf{1})^{-1}A(A_{\varepsilon}+x\mathbf{1})^{-1}dx,\quad G(\varepsilon)=\int_{0}^{\infty}A(B_{\varepsilon}+x\mathbf{1})^{-1}B(A_{\varepsilon}+x\mathbf{1})^{-1}dx.$
We claim that
(50) $\lim_{n\to\infty}\|F(1/n)-F(0)\|=0,\quad lim_{n\to\infty}\|G(1/n)-G(0)\|,$
if the condition (a) holds, and $lim_{n\to\infty}\|\big(F(1/n)+G(1/n)\big)-\big(F(0)+G(0)\big)\|=0$ if the condition (b) holds. Assume first the condition (a). Use the arguments of part (1) to deduce
$\|A(B_{\varepsilon}+x\mathbf{1})^{-1}A(A_{\varepsilon}+x\mathbf{1})^{-1}\|\leqslant\|A(B+(x+\varepsilon)\mathbf{1})^{-1}\|\|A(A+x+\varepsilon\mathbf{1})^{-1}\|\leqslant$
$\left(\frac{\alpha\|B\|}{\|B\|+x+\varepsilon}\right)\left(\frac{\|A\|}{\|A\|+x+\varepsilon}\right)$
Hence,
$\|F(\varepsilon)\|\leqslant\int_{0}^{\infty}\|A(B_{\varepsilon}+x\mathbf{1})^{-1}A(A_{\varepsilon}+x\mathbf{1})^{-1}\|dx\leqslant$
$\alpha\|A\|\|B\|\int_{0}^{\infty}\frac{dx}{(\|B\|+x+\varepsilon)(\|A\|+x+\varepsilon)}dx=$
$\alpha\|A\|\|B\|\frac{\log(\|B\|+\varepsilon)-\log(\|A\|+\varepsilon)}{\|B\|-\|A\|}\leqslant\alpha\|A\|\|B\|\frac{\log(\|B\|+1)-\log(\|A\|+1)}{\|B\|-\|A\|}$
The arguments in part (1) yield that $F_{\varepsilon}$ is well defined for $\varepsilon\in(0,1]$. To estimate $G(\varepsilon)$ we use the inequality $B^{2}\leqslant\beta^{2}A^{2}$ to show
$\|A(B_{\varepsilon}+x\mathbf{1})^{-1}B(A_{\varepsilon}+x\mathbf{1})^{-1}\|\leqslant\|A(B+(x+\varepsilon)\mathbf{1})^{-1}\|\|B(A+x+\varepsilon\mathbf{1})^{-1}\|\leqslant$
$\left(\frac{\alpha\|B\|}{\|B\|+x+\varepsilon}\right)\left(\frac{\beta\|A\|}{\|A\|+x+\varepsilon}\right)\Rightarrow$
$\|G(\varepsilon)\|\leqslant\alpha\beta\|A\|\|B\|\frac{\log(\|B\|+\varepsilon)-\log(\|A\|+\varepsilon)}{\|B\|-\|A\|}\leqslant\alpha\beta\|A\|\|B\|\frac{\log(\|B\|+1)-\log(\|A\|+1)}{\|B\|-\|A\|}$
The equalities (50) follow from Lebesgue’s dominated convergence theorem. This proves (2a).
Assume now the condition (b) $(A-B)^{2}\leqslant\beta^{2}A^{2}$. Then
$\|A(B_{\varepsilon}+x\mathbf{1})^{-1}(A-B)(A_{\varepsilon}+x\mathbf{1})^{-1}\|\leqslant\|A(B+(x+\varepsilon)\mathbf{1})^{-1}\|\|(A-B)(A+(x+\varepsilon)\mathbf{1})^{-1}\|\leqslant$
$\left(\frac{\alpha\|B\|}{\|B\|+x+\varepsilon}\right)\|(A+(x+\varepsilon)\mathbf{1})^{-1}(A-B)^{2}(A+(+\varepsilon)x\mathbf{1})^{-1}\|^{1/2}\leqslant$
$\left(\frac{\alpha\|B\|}{\|B\|+x+\varepsilon}\right)\left(\frac{\beta\|A\|}{\|A\|+x+\varepsilon}\right)$
Hence,
$\|F(\varepsilon)-G(\varepsilon)\|\leqslant\int_{0}^{\infty}\|A(B_{\varepsilon}+x\mathbf{1})^{-1}(A-B)(A_{\varepsilon}+x\mathbf{1})^{-1}\|dx\leqslant$
$\alpha\beta\|A\|\|B\|\frac{\log(\|B\|+\varepsilon)-\log(\|A\|+\varepsilon)}{\|B\|-\|A\|}\leqslant\alpha\beta\|A\|\|B\|\frac{\log(\|B\|+1)-\log(\|A\|+1)}{\|B\|-\|A\|}.$
The arguments of the proof of the case (2a) yield (2b). ∎